If the total cost function of a firm is C=1/3Q3+5Q2+30Q+10 and the price under perfect competition is given as birr 6

A.  Find for what values of Q profit will be maximized

B.  Will the firm continue production at that output level? 



1
Expert's answer
2022-02-17T11:10:57-0500

Given the cost function, "C(q)={1\over3}q^3+5q^2+30q+10" and price, "p=6", the competitive firm’s profit maximization problem can be written as:

"max_{q\ge0}=6q-{1\over3}q^3-5q^2-30q-10"

Differentiating profit with respect to "q" gives marginal profit curve:

Marginal Profit "(q)=6\u2212q^2+10q\u221230"

Equating the marginal profit to 0 and solving,

"-q^2-10q-24=0"

Using the quadratic formula to solve,

"q={-b\pm\sqrt{b^2-4ac}\over 2a}" where, "a=-1,b=-10,c=-24"

So,

"q={10\pm\sqrt{100-96}\over -2}"

Therefore, the values of "q" are "-4" and "-6".


To confirm whether this firm will continue production at this output level, we substitute for the values of "q" in,the function,

"6q-{1\over3}q^3-5q^2-30q-10"

Now, when q=-4, this function gives, "{82\over3}"

When q=-6, the above function gives, 26

For both values above, the profits are greater than 0, which indicates price minimum. Therefore, the firm should not continue with this output level.